OpenStudy (anonymous):

what is the definite integral of x^1/2

OpenStudy (ghazi):

what are the limits of integration?

OpenStudy (anonymous):

\[\Large \int_a^b x^{\frac{1}{2}}dx \] You forgot the limits of integration.

OpenStudy (ghazi):

\[\int\limits_{a}^{b} x^n = [x^{n+1}]/ n+1]\] from a to b

OpenStudy (amistre64):

the rule for integration of powers \[x^n\to\ \frac{1}{n+1}x^{n+1}\]

OpenStudy (anonymous):

For definite integral there must be limits..

OpenStudy (amistre64):

just use a and b until further defined :)

OpenStudy (anonymous):

i've got it...thanks!